The management of Firebolt Industries Inc. manufactures gasoline and diesel engines through two production departments, Fabrication and Assembly. Management needs accurate product cost information in order to guide product strategy. Presently, the company uses a single plantwide factory overhead rate for allocating factory overhead to the two products. However, management is considering the multiple production department factory overhead rate method. The following factory overhead was budgeted for Firebolt:

1

Fabrication Department factory overhead

$557,750.00

2

Assembly Department factory overhead

257,550.00

3

Total

$815,300.00

Direct labor hours were estimated as follows:

Fabrication Department 4,850 hours
Assembly Department 5,050
Total 9,900 hours

In addition, the direct labor hours (dlh) used to produce a unit of each product in each department were determined from engineering records, as follows:

Production Departments Gasoline Engine Diesel Engine
Fabrication Department 3.1 dlh 2.1 dlh
Assembly Department 2.1 3.1
Direct labor hours per unit 5.2 dlh 5.2 dlh
Required:
a. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the single plantwide factory overhead rate method, using direct labor hours as the activity base.*
b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department.*
c. (1) Recommend to management a product costing approach, based on your analyses in (a) and (b). (2) Give a reason for your answer.

*If required, round all per-unit answers to the nearest cent.

b. Determine the per-unit factory overhead allocated to the gasoline and diesel engines under the multiple production department factory overhead rate method, using direct labor hours as the activity base for each department. If required, round all per-unit answers to the nearest cent.

Gasoline engine per unit
Diesel engine per unit
Answer #1

Part (a)

Total factory overhead = $815,300

Total Direct labor hrs = 9900 hrs

Thus, Single Plantwide Overhead rate per direct labor hour = 815300/ 9900 = $ 82.35 per hour

Number of labor hours required per unit = 5.2 hours

Hence, factory overhead allocated per units:

Gasoline Engine = $ 82.35 * 5.2 = $428.22 per unit

Diesel Engine = $ 82.35 * 5.2 = $428.22 per unit

Part (b)

Fabricating Deptt.

Assembly Deptt.

Factory Overhead 557750 257550
No. of direct labor hours 4850 5050
Overhead rate per hour 115 51

Total overhead allocated per unit:

Gasoline Engine = (115 * 3.1) + (51 * 2.1) = $463.60

Diesel Engine = (115*2.1) + (51 * 3.1) = $ 399.60

Part (c)

Product costing approach used under Part (b) should be followed by the management.

Reason: Even though both products use the same number of total direct labor hours, the number of hours consumed in each department is different. Therefore, it is NOT recommended to use a single plant wide overhead rate. Each department should have its own overhead allocation rate. Thus, product costing approach under part (b) should be followed.